3) A box of 25 kg is pushed horizontally, with a force of 200 N, up a 37degrees
(with respect to the horizontal) frictionless incline.
a) Find the Normal force
What is the acceleration of the box?
b) Now if friction is present. What value of the coefficient of kinetic
friction is necessary such that the box is pushed with constant velocity.
Answer #1

a) N + 200sin37 = 25g

N = 25 x 9.81 - 200sin37 =124.89 N

Fcos37 = ma

a = 200cos37 / 25 =6.34 m/s2


b) constant velocity means zero acc.

so friction = Fcos37

u .124.89 = 200cos37

u =1.28